OpenStudy (readergirl12):

Which pairs are solutions to the equation? 5xy + 9 = 44 A. (4, 11) and (2, 3) B. (7, 1) and (2, 3) C. (1, 7) and (4, 11) D. (1, 7) and (7, 1)

pooja195 (pooja195):

Plug in your answer choices into the equation : ) A. 5(4)(11) + 9 = 44 true or false? 5(2)(3) + 9 = 44 true or false?

OpenStudy (readergirl12):

5(4)(11)+9=44 False? 5(2)(3)+9=44 False?

pooja195 (pooja195):

Correct so we know its not A B. We can elminate B because the 2nd coordinate is (2,3) we know its false

pooja195 (pooja195):

Let's try C 5(1)(7) + 9 = 44 <--t or f? 5(4)(11) + 9 = 44 <---t or f?

OpenStudy (readergirl12):

5(1)(7)+9=44 True? 5(4)(11)+9=44 False?

pooja195 (pooja195):

Correct so we know its D We know that (1,7) is correct that leaves us with the 2nd coordinate lets check 5(7)(1)+ 9 = 44 t or f

OpenStudy (readergirl12):

True.

pooja195 (pooja195):

Good so we know D is the answer! =)

OpenStudy (readergirl12):

Thankssss cx

pooja195 (pooja195):

np =P